OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 5 and y-intercept of -3.@Savannah_Noelle

OpenStudy (johnweldon1993):

Slope-intercept form...is written as \[y = mx + b\] 'm' stands for the slope and 'b' stands for the y-intercept So since you were given both slope and the y-intercept...what would the equation look like?

OpenStudy (anonymous):

idk

OpenStudy (johnweldon1993):

So you have....slope of line = 5 y-intercept of line = -3 When applying that to the equation posted above where 'm' is where you put the slope....and 'b' is where you put the y-intercept you would have y = (slope)x + (y-intercept) plug in the numbers you were given

OpenStudy (anonymous):

i dont know how to do it just try this one for me

OpenStudy (anonymous):

y=5x+-3 right or wrong

OpenStudy (johnweldon1993):

after substituting in the values you were given...into that equation...you'll have \[y = 5x - 3\]

OpenStudy (johnweldon1993):

You were correct......but it seems to be a pain putting + -3....that can just be written as -3 so y = 5x - 3

OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 3 and contains the point (4, 9).

OpenStudy (johnweldon1993):

so this one is slightly different...but same equation y = mx + b Now, you have a 'y' value, an 'x' value, and a slope 'm' value...plugging all those in we get 9 = 3(4) + b We need to solve for b here. can you see how to solve for 'b'?

OpenStudy (anonymous):

no

OpenStudy (johnweldon1993):

So when we simplify that 9 = 3(4) + b 3 times 4 is 12 so 9 = 12 + b we can subtract 12 from both sides to isolate 'b' on the right side so -3 = b Now we have 'b' To get the equation of the line you want...again, all you use is the slope 'm' (3)...and the y-intercept 'b' (-3) that we have just found y = (slope)x + (y-intercept) plug in the numbers

OpenStudy (anonymous):

y=3x-3

OpenStudy (johnweldon1993):

there you go!

OpenStudy (anonymous):

and for this one Write the equation in slope-intercept form of the line that has a slope of -3 and contains the point (1, 1).

OpenStudy (johnweldon1993):

Same exact thing that we just did with the previous question...can you follow those steps I did above and figure this one out? Same process, different values

OpenStudy (anonymous):

its hard though

OpenStudy (anonymous):

i dont know if this is correct or not y=2x+1

OpenStudy (johnweldon1993):

That's okay, so we'll go through it again you have a slope 'm' (-3) and a y-value 'y' (1) and an x-value 'x' (1) So plugging all those into that same equation y = mx + b you get 1 = -3(1) + b now we simplify that...-3 times 1 is -3 so 1 = -3 + b add 3 to both sides to isolate 'b' and you get 4 = b remember 'b' is your y-intercept so now you have a slope 'm' (-3) and a y-intercept 'b' (4) so your equation will be y = (slope)x + (y-intercept) your equation is?

OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 2 and contains the point (1, 1).